# Continuous Approximations of Projected Dynamical Systems via Control Barrier Functions

### v1

@@ VII Technical results and proofs
where in the second inequality we use that $0\leq h(x)\leq-\frac{1}{\alpha}\mathcal{L}_{f}h(z,x)$. ∎
Lemma VII.2.
For any $\alpha\geq\alpha_{*}$ and $z\in\mathcal{Z}$, for all $x\in U_{\mathrm{cbf},\alpha}(z)$, we have
$0<\epsilon\leq\|\nabla h(x)\|\leq M_{3}$
Proof of Lemma VII.2.
@@
$\|\nabla h(x)-\nabla h(y)\|\leq L_{\nabla h}\|x-y\|\implies$
$\|\nabla h(x)\|\geq\|\nabla h(y)\|-L_{\nabla h}\|x-y\|\implies$
@@
where in the last step we have used Lemma VII.1. Thus, for any $\alpha\geq\alpha_{*}$, and for any $x\in U_{\mathrm{cbf},\alpha}(z)$, we get:
$\|\nabla h(x)\|$ $\geq M_{1}-L_{\nabla h}\gamma(\frac{1}{\alpha}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|)\geq M_{1}-L_{\nabla h}\gamma(\frac{1}{\alpha_{*}}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|)=\epsilon>0$
Similarly, we can deduce that for any $\alpha\geq\alpha_{*}$, and for any $x\in U_{\mathrm{cbf},\alpha}(z)$:
$\|\nabla h(x)-\nabla h(y)\|\leq L_{\nabla h}\|x-y\|\implies$
[…]
Lemma VII.3.
@@
which is well-defined, as $\nabla h(x)\neq 0$, from Lemma VII.2.
@@
where in the second inequality we used Assumption 2, in the third inequality we used Lemma VII.3, in the fourth inequality we used that $|\mathcal{L}_{f}h(z,x)+\alpha h(x)|\leq|\mathcal{L}_{f}h(z,x)|$, and in the fifth inequality we used Lemma VII.1.
@@
where in the fifth inequality we used $\|\nabla h(x)\|\leq\|\nabla h(y)\|+L_{\nabla h}\|x-y\|$, as proven in the proof of Lemma VII.2, and in the last inequality we used Lemma VII.1.
@@
The proof follows steps similar to [10, Section 4]. In what follows, we prove graphical convergence [^7] For the definition of graphical convergence, see [25, Def. 5.18]. of $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ or $\{\phi_{i}\}_{i\in\mathbb{N}}$ to $\phi$. Thus, we regularly treat $\phi_{i_{k}}$, $\phi_{i}$ and $\phi$ as sets, through their graphs, below. From [25, Lemma 5.28], it follows that graphical convergence implies uniform convergence on compact subintervals of $[0,T)$. Further, note that, since $\sigma(\cdot)$ is decreasing and $\lim_{\alpha\to\infty}\sigma(\alpha)=0$, the sequence $\sigma(\alpha_{i})$ is decreasing and $\lim_{i\to\infty}\sigma(\alpha_{i})=0$.
Proof of Statement 1. From [25, Thm. 5.7], $\{\phi_{i}\}_{i\in\mathbb{N}}$ has a graphically convergent subsequence, say $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$. Note that $\phi_{i}$ are also solutions to perturbed DI (11) for $\alpha=\alpha_{i}$ (Cor. IV.1), i.e. $\phi_{i}$, and thus $\phi_{i_{k}}$, are solutions to $\sigma(\alpha_{i})$-perturbations, and $\sigma(\alpha_{i_{k}})$-perturbations, respectively, of the limiting DI (12) (Cor. IV.2). Thus, since the limiting DI (12) is well-posed (Prop. VII.1), and since $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ is convergent, then $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ graphically converges to a solution $\phi$ of the limiting DI (12), from [25, Def. 6.29]. Finally, from Cor. IV.3 item 1, $\phi$ is a solution to PDS (7).
Proof of Statement 2. Follows identical steps to the proof of [10, Cor. 4.6]. ∎
@@
Let $\gamma^{\prime}>0$ with $\gamma^{\prime}<\gamma$. Under our assumptions, the limiting DI (12) is well-posed (Prop. VII.1) and it has the exact same solutions as the PDS (7) (from Cor. IV.3). Moreover, perturbed DI (11) is a $\sigma(\alpha)$-perturbation to limiting DI (12) (Cor. IV.2). Then, from [25, Prop. 6.34], it follows that there exists $\alpha^{\prime}\geq\alpha_{*}$, such that for all $\alpha\geq\alpha^{\prime}$ every solution $\chi$ to perturbed DI (11) is $\gamma^{\prime}$-close to a solution $\phi$ of (12), implying $\chi(t)\in(\mathcal{Z}^{\prime}+\gamma^{\prime}\mathbb{B})\times\mathcal{S}$, for all $t\in[0,T]$.

### v3

@@ VIII Technical results and proofs
where we used $0\leq h(x)\leq-\frac{1}{\alpha}\mathcal{L}_{f}h(z,x)$. ∎
Lemma VIII.2.
For any $\alpha\geq\alpha_{*}$ and $z\in\mathcal{Z}$, for all $x\in U_{\mathrm{cbf},\alpha}(z)$, we have $0<\epsilon\leq\|\nabla h(x)\|\leq M_{3}$.
Proof of Lemma VIII.2.
@@
$\|\nabla h(x)-\nabla h(y)\|\leq L_{\nabla h}\|x-y\|\implies\|\nabla h(x)\|\geq\|\nabla h(y)\|-L_{\nabla h}\|x-y\|\implies$
@@
where in the last step we have used Lemma VIII.1. Thus, for any $\alpha\geq\alpha_{*}$, and for any $x\in U_{\mathrm{cbf},\alpha}(z)$, we get:
$\|\nabla h(x)\|$ $\geq M_{1}-L_{\nabla h}\gamma(\frac{1}{\alpha}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|)$
$\geq M_{1}-L_{\nabla h}\gamma(\frac{1}{\alpha_{*}}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|)=\epsilon>0$
Similarly, for any $\alpha\geq\alpha_{*}$ and for any $x\in U_{\mathrm{cbf},\alpha}(z)$, $\|\nabla h(x)-\nabla h(y)\|\leq L_{\nabla h}\|x-y\|$ implies:
$\|\nabla h(x)\|\leq\|\nabla h(y)\|+L_{\nabla h}\|x-y\|$ $\leq M_{2}+L_{\nabla h}\gamma(\frac{1}{\alpha}|\mathcal{L}_{f}h(z,x)|)$
$\leq M_{2}+\gamma(\frac{1}{\alpha_{*}}|\mathcal{L}_{f}h(z,x)|)$
$\leq M_{3}$
Lemma VIII.3.
@@
which is well-defined, as $\nabla h(x)\neq 0$, from Lemma VIII.2.
@@
where in the second inequality we used Assumption 2, in the third one we used Lemma VIII.3, in the fourth one we used $|\mathcal{L}_{f}h(z,x)+\alpha h(x)|\leq|\mathcal{L}_{f}h(z,x)|$, and in the fifth one we used Lemma VIII.1.
@@
where in the fifth inequality we used $\|\nabla h(x)\|\leq\|\nabla h(y)\|+L_{\nabla h}\|x-y\|$, as proven in the proof of Lemma VIII.2, and in the last one we used Lemma VIII.1.
@@
We follow steps similar to [10, Section 4]. We prove graphical convergence [^7] For the definition of graphical convergence, see [25, Def. 5.18]. of $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ or $\{\phi_{i}\}_{i\in\mathbb{N}}$ to $\phi$. Thus, we regularly treat $\phi_{i_{k}}$, $\phi_{i}$ and $\phi$ as sets, through their graphs, below. From [25, Lemma 5.28], graphical convergence implies uniform convergence on compact subintervals of $[0,T)$. Further, since $\sigma(\cdot)$ is decreasing and $\lim_{\alpha\to\infty}\sigma(\alpha)=0$, the sequence $\sigma(\alpha_{i})$ is decreasing and $\lim_{i\to\infty}\sigma(\alpha_{i})=0$.
Proof of Statement 1. From [25, Thm. 5.7], $\{\phi_{i}\}_{i\in\mathbb{N}}$ has a graphically convergent subsequence, say $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$. Note that $\phi_{i}$ are also solutions to perturbed DI (11) for $\alpha=\alpha_{i}$ (Cor. IV.1), i.e., $\phi_{i}$, and thus $\phi_{i_{k}}$, are solutions to $\sigma(\alpha_{i})$-perturbations, and $\sigma(\alpha_{i_{k}})$-perturbations, respectively, of the limiting DI (12) (Cor. IV.2). Thus, since the limiting DI (12) is well-posed (Prop. VIII.1), and since $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ is convergent, then $\{\phi_{i_{k}}\}_{k\in\mathbb{N}}$ graphically converges to a solution $\phi$ of the limiting DI (12), from [25, Def. 6.29]. Finally, from Cor. IV.3 item 1, $\phi$ is a solution to PDS (7).
Proof of Statement 2. Similar to the proof of [10, Cor. 4.6]. ∎
@@
Let $\gamma^{\prime}>0$ with $\gamma^{\prime}<\gamma$. Limiting DI (12) is well-posed (Prop. VIII.1) and has the same solutions as PDS (7) (Cor. IV.3). Perturbed DI (11) is a $\sigma(\alpha)$-perturbation to limiting DI (12) (Cor. IV.2). Then, from [25, Prop. 6.34], $\exists\alpha^{\prime}\geq\alpha_{*}$, such that for all $\alpha\geq\alpha^{\prime}$ every solution $\chi$ to perturbed DI (11) is $\gamma^{\prime}$-close to a solution of (12), implying $\chi(t)\in(\mathcal{Z}^{\prime}+\gamma^{\prime}\mathbb{B})\times\mathcal{S}$, $\forall t\in[0,T]$.